Suppose that the utility function for two

commodities is given by U = x2 yand the budget constraint is 10x + 15y = 250. Find the values of x and y that maximize utility 



1
Expert's answer
2022-01-31T09:34:14-0500

The utility is maximized, when:

MUx/Px = MUy/Py and 10x + 15y = 250.

"MUx = U'(x) = 2xy^3,"

"MUy = U'(y) = 3x^2 y^2."

"2xy^3\/10 = 3x^2 y^2\/15,"

y/5 = x/5,

y = x,

10x + 15x = 250,

x = y = 10 units.